Use the following information to answer Questions 3 to 5: An analyst gathered the following information regarding Pluto Inc: •Expected EPS for 2011 = $4.25 •Retention rate = 0.4 •Required return = 11% •Current stock price = $54.85 Dividends are paid out at the end of the year and are expected to grow at a rate of 5.5% forever.

Question: The intrinsic value of the stock at the end of 2010 is closest to:

Select one: a. $46.36 b. $77.27 c. $30.91

Question text The present value of growth opportunities is closest to: Select one: a. $31.67 b. $16.21 c. $7.72

Question text The fraction of the company's leading P/E ratio that comes from present value of growth opportunities is closest to: Select one: a. 3.81% b. 29.55% c. 34.97%
Answer #1

1

Dividend = EPS*(1-retention ratio)

=4.25*(1-0.4) = 2.55

As per DDM
Price = Dividend in 1 year* (1 + growth rate )/(cost of equity - growth rate)
Price = 2.55/ (0.11 - 0.055)
Price = 46.36

2

PVGO = intrinsic value - EPS/cost of equity = 46.36-4.25/.11=7.72

3

Fraction of PVGO in leading P/E

=(current price-EPS/cost of equity)/current price = (54.85-4.25/.11)/54.85 = 29.55%